# Grounded in Literature

### v1

@@ 3.2 Evaluation Protocol
A high-quality research idea evaluation should highlight all strengths and weaknesses that an expert reviewer would mention, ground its claims in relevant literature, engage deeply with the points discussed and the works cited, and go beyond simple good or bad judgment to offer actionable suggestions for improvement. Guided by this desiderata, we develop automatic metrics to assess the performance of ScholarEval and baselines. Full details of prompts and implementation are provided in the Appendix D.
@@
. This is a recall-based metric which measures the extent to which an evaluation covers the rubrics from the corresponding review in ScholarIdeas. Specifically, we use Prometheus-Eval with a GPT-4 backbone (OpenAI et al., 2024), a framework shown to follow detailed evaluation rubrics effectively (Kim et al., 2024). We instruct it to assign a 1–5 score based on how well each evaluation addresses the points in the reference rubric. The final score is the average over all 1,076 rubrics in ScholarIdeas.
@@
, Act, Depth. We use an LLM-as-judge, Claude 4 Sonnet, to give pairwise preferences between ScholarEval and our strongest baseline along three criteria: Evidence Support, which measures how well claims are grounded in literature and supported by relevant citations; Actionability, which captures the clarity, usefulness, and feasibility of the suggestions; and Depth, which evaluates the level of engagement with each point and whether the evaluation mentions specifics about the work it cites rather than relying on generic statements. We also compute agreement with human annotators (see Appendix D).
@@ 4.1 Experimental Details
Baselines. Because there are no existing systems in the literature that evaluate research ideas with feedback at the same scope as ScholarEval, we compare against strong baselines that are most likely to be used for idea evaluation and feedback. Specifically, we select both frontier open-source and closed-source LLMs—Llama-3.3-70B (Grattafiori et al., 2024), GPT-4.1 (OpenAI, 2025a), Claude-4-Sonnet (Anthropic, 2025), and GPT-4o-search-preview (OpenAI, 2025c) for a web-connected LLM baseline—as well as a deep research system, o4-mini-deep-research (OpenAI, 2025b). We restrict our choice of deep research systems o4-mini-deep-research since it is available via API. All baselines are prompted with a detailed template aligned with ScholarEval, including method and dimension decomposition and dedicated sections for strengths, weaknesses, and suggestions.
@@
Full prompts and additional experimental setup details are provided in Appendix E.
@@ 4.2 Results and Analysis
Table 1: Coverage $\uparrow$ of baselines and variants of ScholarEval overall and per-discipline. and indicate significant improvement over 1 or all baselines, respectively. Best results are bolded. Statistical significance details are provided in Appendix F.
Overall AI Neuroscience Biochemistry Ecology
o4-mini-deep-research 2.28 $\pm$ 1.07 2.35 $\pm$ 1.07 2.25 $\pm$ 1.04 2.08 $\pm$ 1.06 2.35 $\pm$ 1.11
ScholarEval $_{\text{GPT}}$ 2.72 $\pm$ 1.47 2.84 $\pm$ 1.39 2.74 $\pm$ 1.55 2.61 $\pm$ 1.42 2.52 $\pm$ 1.51
Figure 4: Win rate based on Depth, Act, and Evidence between ScholarEval $_{\text{Claude}}$ and o4-mini-deep-research using LLM judge.
@@
o4-mini-deep-research 1.07%
@@
ScholarEval $_{\text{GPT}}$ 0%
@@
[…]
ScholarEval eliminates reference invalidity. As shown in subsection 4.2, only ScholarEval variants achieve 0% Reference Inv. $\downarrow$ in our evaluation. As previously stated, baseline rates represent lower bounds. However, a manual audit of a sample of outputs (see §F.2) indicates that up to 80% of Llama-3.3-70B citations are hallucinated in certain evaluations. Even GPT-4o-search-preview and o4-mini-deep-research are not immune, with at least 1% of cited references not existing. Our audit also uncovers subtler failures, including misattributed authors and claims that are inconsistent with the cited sources. These patterns align with findings from recent evaluations of deep research systems (Li et al., 2025a). ScholarEval eliminates these issues, ensuring that all citations resolve to valid, traceable sources.
@@
[…]
Ablation studies. We conduct ablations to assess the effectiveness of individual components of ScholarEval. Namely, we remove each of the methods and results extraction (MRE), paper augmentation (PA), and pairwise comparison (PC). Details about the setup for each ablation experiment can be found in Appendix E. Results in Table 3 indicate that the removal of each of these components leads to degradation in performance of ScholarEval on Coverage $\uparrow$. This is likely due to the reduced information given to the model (i.e. the effectiveness of the methods from prior work based on the reported results, the dense list of relevant papers, and the granular dimension-level paper comparisons). These results showcase the importance of these steps in providing ScholarEval with essential context from the literature to generate effective idea evaluations.
@@ 5 Expert User Study
Design. We gauge the real-world usefulness of ScholarEval by recruiting 18 experts for 46 total evaluations in a blind experiment with OpenAI’s o4-mini-deep-research OpenAI (2025b). Each expert had an education level of PhD student or beyond and was verified to have at least one paper published in their field (see Table 16). We create a blind interface for experts to interact with ScholarEval (see Figure 6 and Figure 7).
@@
Procedure. We ask each expert participant to first write a research idea that includes the problem they aim to tackle and the suggested methodology. This can be one they have yet to experiment with, or an idea they have already published. After receiving the soundness and contribution evaluation from a single system (either ScholarEval or OpenAI deep research), they complete a detailed rubric related to the core components of idea evaluation (e.g. the usefulness of the suggestions, the engagement with literature, etc.). The entire process for creating an idea, generating the feedback, and scoring the rubric took an average of 1 hour across participants, and we allow up to 4 unique idea submissions. The exact questions, user demographics, blindness and randomization validity, and compensation details can be found in Appendix G.
@@
Evaluation. We group questions into six dimensions. Citations captures the number of references experts would actually use in their research. Because research ideas can be complex and domain-specific, we use Faithful to measure the extent to which the feedback aligns with nuances of the research idea. Useful reflects the helpfulness of each system and the expert’s enthusiasm for future use. Focus measures the degree to which each system targets the most important factors of each research idea. LitEngage indicates the depth each system uses when making detailed comparisons with specific components of relevant literature. Refine gauges whether the system provides valuable, targeted, feasible suggestions for improvement that experts believe would actually improve their research idea. The questions related to each dimension are available in Table 15. We refer readers to §G.2 for details on linear mixed effects model, which we adopt as our statistical method.
Table 4: Mixed effects modeling results from the expert user study. We report the standardized regression coefficient ($\beta$) and statistical significance (* $p<.05$, *** $p<.001$). Deep Research refers to o4-mini-deep-research.
Dimension $\beta$ ScholarEval Deep Research
@@
Results and analysis. Results in Table 4 show a statistically significant preference for ScholarEval over o4-mini-deep-research across all six dimensions measured in our expert user study. Experts found 1.5 more useful Citations using ScholarEval and scored LitEngage 1.2 higher, demonstrating the effectiveness of our multi-stage literature retrieval pipeline. A strong effect on Faithful and Focus underscores the benefit of systemically breaking down the research plan into smaller, controlled comparisons with relevant literature. The actionable feedback generated by ScholarEval had clear advantages as well, with experts reporting a 1.2 increase in Refine. The higher score on Useful further underscores the overall usefulness of ScholarEval as research evaluation framework as judged by experts.
6 Related Work
Literature Grounded Systems for Research. Multiple works have been proposed that use LLMs in literature grounded systems to assist in research. A commonly targeted use case for these systems is literature synthesis and related work generation (Asai et al., 2024; Kang et al., 2023; Agarwal et al., 2025). In addition, other systems have been developed for literature understanding and question answering (Skarlinski et al., 2024; Singh et al., 2025). ScholarEval builds on techniques inspired by these works but focuses on the underexplored problem of literature-grounded research idea evaluation.
Research Ideation. Recently, there has been growing interest in using LLMs for generating research ideas and scientific hypotheses, often from a literature-driven perspective. For example, Wang et al. (2024) proposed SciMON, a framework that retrieves inspirations from past scientific papers to generate novel ideas. Similarly, Baek et al. (2025) employ an iterative approach over related papers and a knowledge store to generate scientific hypotheses. Other works, such as (Radensky et al., 2025; Garikaparthi et al., 2025; Pu et al., 2025), emphasize human-LLM collaboration for idea generation, while Gottweis et al. (2025) use a multi-agent approach to generate and refine hypotheses. Furthermore, Si et al. (2025b) conducted a large-scale study assessing both human- and LLM-generated ideas in the AI field, finding that LLMs are capable of generating more novel ideas based on human evaluation.
@@ 7 Conclusion
Limitations of ScholarEval. First, as a literature-grounded framework, ScholarEval relies heavily on the retrieved literature to evaluate research ideas. Hence, it is possible that it might misjudge a method’s effectiveness if it is not yet proven in the existing literature. Similarly, the contribution of a research idea might be misrepresented if some relevant papers are not retrieved for comparison. A meaningful direction for future work would be to improve the retrieval capabilities of ScholarEval to reduce such occurrences. Second, as we mention in Appendix F, depending on the choice of the model backbone, running an idea evaluation using ScholarEval can take around 12 min and cost up to \$3. Although not too high in absolute terms, these costs can add up if ScholarEval is used to evaluate a large batch of ideas. Future work may explore more lightweight versions of ScholarEval to enhance its usability in such cases.
@@
Limitations of Expert User Study. Some measures, such as those related to Focus, are inherently subjective to the user, and could be influenced by stylistic factors rather than actual improved quality. Additionally, resarch idea quality is a source of variance, as participants were only allowed to submit unique research ideas to a given system. However, we collect a self-report of research idea detail, which is not statistically significantly different according to a Welch’s t-test. Additionally, the overall scores of our results may be inflated by our expert sample, who report a 7.2/10 on AI use for research. Our results may be different in a sample who use AI less often. This does not affect our relative improvement over o4-mini-deep-research, as we show concrete evidence that the study was adequately blinded (see Appendix G). Future work should consider the barriers to adoption for automated research idea evaluation, especially for those who are less keen on using AI for their research cycle. Another strong direction would study how use of ScholarEval impacts short and long-term research success.
Reproducibility Statement
To ensure the reproducibility of our work, we provide full prompts and implementation details of ScholarEval (Appendix A), the detailed process of constructing ScholarIdeas (Appendix C), full information about our evaluation (Appendix D), experimental setup (Appendix E), and expert user study (Appendix G). We also openly release our code and dataset.
@@
ScholarIdeas is collected from papers on Openreview and eLife licensed under a Creative Commons Attribution (CC BY) license, which allows reuse with attribution. We provide the full list of papers used to create ScholarIdeas in Appendix C.
@@ Acknowledgments
Appendix A ScholarEval details and prompts
@@
A.1 Soundness Details
A.1.1 Additional Implementation Details
@@
TL;DR Summary. Since the method-level soundness review can be lengthy, we also include a summarization step where we prompt an LLM to generate a TL;DR summary highlighting the top three most important strengths, weaknesses, and suggestions to address. In our ScholarEval user interface (Appendix B), both this summary and the method-level evaluation are shown to the user, with the latter being expandable.
@@ A.1.2 Prompts
Methods Extraction You are an expert research assistant. You are skilled at reading research ideas and identifying the methods that are being proposed to solve the research problem. Methods can be planned system designs, experiments, human studies, analyses, ablations, etc. Given a research idea, you should extract all methods as a Python list, such that each method is a separate item in the list. Each item should be a word-for-word copy of a method, along with a short synthesis that grounds the method in the context of the overall research idea. The extracted methods should be interpretable on their own. Ensure that the methods you extract address different aspects of the research idea and are non-redundant. The method list you return should be ranked by importance to the research idea, with the most important methods first. [start research idea] {research_idea} [end research idea] Please output a parseable Python block as follows:
‘‘‘python
plans = ["context + method", ...]
@@
Query Generation You are an expert research assistant. Given a method (i.e., one approach that researchers are adopting to execute their idea) extracted from a research idea, please construct a singular query that will be used to search for paper snippets using the Semantic Scholar API. Use JSON format with 70 words or less per query. Do not include any text in the query besides the query itself. Do not include text like "semantic search query about …" or "papers related to …". Just the actual query text. No operators such as AND, OR should be used. Just a query in natural language that is relevant to the method. [start extracted method] {clean_method} [end extracted method] In case the method does not have enough context to construct an effective snippet search query, you can use the research idea to understand the overall research direction and inject useful context. [start research idea] {research_idea} [end research idea] Please output a parseable JSON block as follows, being especially careful to use the correct number of escape characters:
‘‘‘json
    "query": "Your search query here (IN 70 WORDS OR LESS)"
@@
Method and Result Synthesis You are an expert research assistant knowledgeable in many domains. You are extremely critical and observant, and do not overgeneralize findings. You are given a proposed research method and the methods/results section from a paper. [start paper] {paper_text} [end paper] [start proposed research method] {proposed_method} [end proposed research method] To further understand the scope of the proposed research method, here is the entire research idea that it is extracted from — a method is a single approach that researchers are adopting to execute their research idea: [start research idea] {research_idea} [end research idea] For any method in the paper that is related to the proposed research method and the overall research idea, please summarize the method used in the paper, report the experimental outcome from using the method, and provide some context for experimental conditions. Do not use any in-text citations. Ensure that the method, results, and context you provide are specific and detailed, and that they mention how it relates to the proposed method and research idea. If the proposed research method does not relate to any methods in the paper, please return an empty dictionary. Strictly follow the output format displayed below. JSON formatting requirements: - Must be a complete, valid JSON object - Start with an open bracket and end with closed bracket - No trailing commas after the last property - Validate JSON structure before output
‘‘‘json
    "method": "Description of experimental approach including:
    algorithm/technique, datasets/inputs, computational
    resources, implementation/experimentation details, and
    evaluation setup, and metrics/instruments used, etc.",
    "results": "Quantitative outcomes with specific values,
    comparisons to baselines, statistical significance where
    applicable",
    "context": "Key experimental conditions:
    dataset/population size, hardware/system/instrument
    specs, hyperparameters, or other domain-specific
    constraints that affect reproducibility"
@@
Method-Level Soundness Review Synthesis You are an expert research assistant knowledgeable in many domains. You are extremely critical and observant, and do not overgeneralize findings. You are given a proposed research method and a list of related work. Your objective is to create a meta-review of the related work in the context of the proposed research method. Point out any evidence that supports or contradicts the proposed method. Make sure to contrast the related work as a series of iterative scientific work, where newer work can provide evidence that supports or contradicts older work. It is important that the meta-review you generate always ties back to the original research idea. Judge the support and contradictions as well as suggested actions for each method within the general context of the research idea to ensure that your review is highly relevant and precise. Be granular, making sure to reference specific details such as: algorithm/technique, datasets/inputs/population, computational resources, statistical methods, implementation details, and evaluation setup, and metrics/instruments used, etc. quantitative outcomes, comparisons to baselines, statistical significance dataset/population size, hardware/system specs, hyperparameters, or other domain-specific constraints that affect reproducibility It is important that for each method-level meta-review that you generate, your review of the support and contradictions should be ordered starting from strongest evidence of support/contradiction to the weakest. Likewise, the suggested actions should be ordered from most important to least important. This does not mean that you will generate these as bullet points, but rather detailed, coherent paragraphs that are logically ordered. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Related Work: {related_work} [start proposed research method] {pm} [end proposed research method] [start research idea] {rp} [end research idea] Output format: Please output a parseable JSON block as follows:
‘‘‘json
    "support": "evidence that supports the proposed method",
    "contradictions": "evidence that contradicts the proposed
                       method",
    "suggested_action": "how can the proposed method be improved
                        based on the related work",
    "soundness_score": "int score 0 to 10 based on the evidence
                        for and against the proposed method"
A.2 Contribution Details
A.2.1 Additional Implementation Details

### v2

@@ 3.2 Evaluation Protocol
Overall AI Neuroscience Biochemistry Ecology
Language Models
GPT-5.1 Instant 1.94 $\pm$ 0.98 2.09 $\pm$ 0.99 1.85 $\pm$ 0.99 1.68 $\pm$ 0.91 1.95 $\pm$ 0.97
Web-connected LM
Deep Research
OpenAI Deep Research 2.28 $\pm$ 1.07 2.35 $\pm$ 1.07 2.25 $\pm$ 1.04 2.08 $\pm$ 1.06 2.35 $\pm$ 1.11
DR Tulu 2.33 $\pm$ 1.14 2.35 $\pm$ 1.11 2.31 $\pm$ 1.17 2.24 $\pm$ 1.04 2.39 $\pm$ 1.20
Ours
ScholarEval $_{\text{GPT-4.1}}$ 2.72 $\pm$ 1.47 2.84 $\pm$ 1.39 2.74 $\pm$ 1.55 2.61 $\pm$ 1.42 2.52 $\pm$ 1.51
ScholarEval $_{\text{GPT-5.1}}$ 2.17 $\pm$ 1.27 2.31 $\pm$ 1.30 2.12 $\pm$ 1.24 1.96 $\pm$ 1.19 2.09 $\pm$ 1.28
Table 1: Coverage $\uparrow$ of baselines and variants of ScholarEval overall and per-discipline. indicates significant improvement over at least one baseline and indicates significant improvement over all baselines. Best results are in bold. Statistical significance details are provided in Appendix G.
A high-quality research idea evaluation should highlight all strengths and weaknesses that an expert reviewer would mention, ground its claims in relevant literature, engage deeply with the points discussed and the works cited, and go beyond simple good or bad judgment to offer actionable suggestions for improvement. Guided by this desiderata, we develop automatic metrics to assess the performance of ScholarEval and baselines. Full details of prompts and implementation are provided in the Appendix Appendix E.
@@
. This is a recall-based metric which measures the extent to which an evaluation covers the rubrics from the corresponding review in ScholarIdeas. Specifically, we use Prometheus-Eval (Kim et al., 2024) with a GPT-4 backbone (OpenAI et al., 2024), a framework that is widely used to evaluate long-form output and that is shown to align well with human judgment (Asai et al., 2024; Yifei et al., 2025; Shao et al., 2024; Bonomo et al., 2025). We instruct it to assign a 1–5 score based on how well each evaluation addresses the points in the reference rubric. The final score is the average over all 1,076 rubrics in ScholarIdeas.
@@
, Act, Depth. We use an LLM-as-judge, Claude 4 Sonnet, to give pairwise preferences between ScholarEval and our strongest baseline along three criteria: Evidence Support, which measures how well claims are grounded in literature and supported by relevant citations; Actionability, which captures the clarity, usefulness, and feasibility of the suggestions; and Depth, which evaluates the level of engagement with each point and whether the evaluation mentions specifics about the work it cites rather than relying on generic statements. We also compute agreement with human annotators, highlighting strong alignment between the LLM judge and human judgment (see Appendix E).
@@ 4.1 Experimental Details
Baselines. We compare ScholarEval against baselines that are capable of providing detailed research idea feedback at a similar scope. Specifically, we select a host of strong baselines consisting of frontier open-source and closed-source LLMs—Llama-3.3-70B (Grattafiori et al., 2024), GPT-4.1 (OpenAI, 2025d), GPT-5.1 Instant (OpenAI, 2025b), Claude-4-Sonnet (Anthropic, 2025), and GPT-4o-search-preview (OpenAI, 2025a) as a web-connected LLM baseline—in addition to two deep research systems: OpenAI Deep Research444o4-mini-deep-research (OpenAI, 2025c), a strong and widely used deep research system and the only proprietary one available via API at the time of writing, and DR-Tulu (Shao et al., 2025b), a recent open-source deep research system that rivals proprietary ones and shows strong performance for academic tasks.
While there exist related academic systems for research idea evaluation, these primarily focus on idea scoring or ranking (Feng et al., 2025; Wen et al., 2025), or provide only short rationales (Shahid et al., 2025), and thus do not offer feedback at the same scope as ScholarEval. We therefore do not treat them as primary baselines, but provide a small-scale comparison in Appendix G for context.
We prompt baselines with a detailed template aligned with ScholarEval output format and depth, including method and dimension decomposition and dedicated sections for strengths, weaknesses, and suggestions.
@@
Full prompts and additional experimental setup details are provided in Appendix F.
@@ 4.2 Results and Analysis
[…]
ScholarEval eliminates reference invalidity. As shown in Table 2, ScholarEval variants and DR-Tulu achieve 0% Reference Inv. $\downarrow$ in our evaluation, while OpenAI Deep Research exhibits a similarly low invalidity rate of 1%, reflecting that information attribution has become increasingly robust in modern research-oriented systems. As previously stated, baseline rates represent lower bounds. However, a manual audit of a sample of outputs (see subsection G.2) indicates that up to 80% of Llama-3.3-70B citations are hallucinated in certain evaluations. Our audit also uncovers subtler failures, including misattributed authors and claims that are inconsistent with the cited sources. These patterns align with findings from recent evaluations of deep research systems (Li et al., 2025a). In contrast, ScholarEval eliminates these issues entirely, ensuring that all citations resolve to valid, traceable sources.
[…]
@@
OpenAI Deep Research 1.07%
DR Tulu 0%
@@
ScholarEval $_{\text{GPT-4.1}}$ 0%
ScholarEval $_{\text{GPT-5.1}}$ 0%
@@
Metric vs. OpenAI DR vs. DR-Tulu
Evidence 76.6% 64.0%
Act 71.2% 62.2%
Depth 83.8% 66.7%
Table 3: Pairwise win rate between ScholarEval $_{\text{Claude}}$ (our best performing system) and OpenAI Deep Research and DR Tulu using LLM judge.
Ablations studies. We conduct ablations to assess the effectiveness of individual components of ScholarEval. Namely, we remove each of the methods and results extraction (MRE), paper augmentation (PA), and pairwise comparison (PC). Details about the setup for each ablation experiment can be found in Appendix F. Results in Table 4 indicate that the removal of each of these components leads to degradation in performance of ScholarEval on Coverage $\uparrow$. This is likely due to the reduced information given to the model (i.e. the effectiveness of the methods from prior work based on the reported results, the dense list of relevant papers, and the granular dimension-level paper comparisons). These results showcase the importance of these steps in providing ScholarEval with essential context from the literature to generate effective idea evaluations.
@@ 5 Expert User Study
Design. We bolster our automatic evaluation with a large-scale blinded user study to further gauge the real-world usefulness of ScholarEval. We recruit 18 experts for 46 total evaluations in a blind experiment with OpenAI Deep Research OpenAI (2025c). Each expert had an education level of PhD student or beyond and was verified to have at least one paper published in their field (see Table 19). We create a blind interface for experts to interact with ScholarEval or OpenAI Deep Research (see Figure 5 and Figure 6). We choose OpenAI DeepResearch as it was both the strongest and most popular model for our task at the time of conducting this experiment.
@@
Procedure. We ask each expert participant to first write a research idea that includes the problem they aim to tackle and the suggested methodology. This can be one they have yet to experiment with, or an idea they have already published. After receiving the soundness and contribution evaluation from a single system (either ScholarEval or OpenAI deep research), they complete a detailed rubric related to the core components of idea evaluation (e.g. the usefulness of the suggestions, the engagement with literature, etc.). The entire process for creating an idea, generating the feedback, and scoring the rubric took an average of 1 hour across participants, and we allow up to 4 unique idea submissions. The exact questions, user demographics, blindness and randomization validity, and compensation details can be found in Appendix H.
@@
Evaluation. We group questions into six dimensions. Citations captures the number of references experts would actually use in their research. Because research ideas can be complex and domain-specific, we use Faithful to measure the extent to which the feedback aligns with nuances of the research idea. Useful reflects the helpfulness of each system and the expert’s enthusiasm for future use. Focus measures the degree to which each system targets the most important factors of each research idea. LitEngage indicates the depth each system uses when making detailed comparisons with specific components of relevant literature. Refine gauges whether the system provides valuable, targeted, feasible suggestions for improvement that experts believe would actually improve their research idea. The questions related to each dimension are available in Table 18. We refer readers to H.2 for details on linear mixed effects model, which we adopt as our statistical method.
Results and analysis. Results in Table 5 show a statistically significant preference for ScholarEval over OpenAI Deep Research across all six dimensions measured in our expert user study. Experts found 1.5 more useful Citations using ScholarEval and scored LitEngage 1.2 higher, demonstrating the effectiveness of our multi-stage literature retrieval pipeline. A strong effect on Faithful and Focus underscores the benefit of systemically breaking down the research plan into smaller, controlled comparisons with relevant literature. The actionable feedback generated by ScholarEval had clear advantages as well, with experts reporting a 1.2 increase in Refine. The higher score on Useful further underscores the overall usefulness of ScholarEval as research evaluation framework as judged by experts.
6 Related Work
Literature Grounded Systems for Research. Multiple works have been proposed that use LLMs in literature grounded systems to assist in research. A commonly targeted use case for these systems is literature synthesis and related work generation (Asai et al., 2024; Kang et al., 2023; Agarwal et al., 2025; Ding et al., 2025). In addition, other systems have been developed for literature understanding and question answering (Skarlinski et al., 2024; Singh et al., 2025; Volkova et al., 2025). ScholarEval builds on techniques inspired by these works but focuses on the underexplored problem of literature-grounded research idea evaluation.
Research Ideation. Recently, there has been growing interest in using LLMs for generating research ideas and scientific hypotheses, often from a literature-driven perspective. For example, Wang et al. (2024) proposed SciMON, a framework that retrieves inspirations from past scientific papers to generate novel ideas. Similarly, Baek et al. (2025) employ an iterative approach over related papers and a knowledge store to generate scientific hypotheses. Other works emphasize human-LLM collaboration for idea generation (Radensky et al., 2025; Garikaparthi et al., 2025; Pu et al., 2025; Shao et al., 2025a; Liu et al., 2025), while multi-agent approaches to generate and refine hypotheses increase in popularity (Gottweis et al., 2025; Ueda et al., 2025; Mitchener et al., 2025). Furthermore, Si et al. (2025b) conducted a large-scale study assessing both human- and LLM-generated ideas in the AI field, finding that LLMs are capable of generating more novel ideas based on human evaluation.
Dimension $\beta$ ScholarEval OpenAI DR
@@ 6 Related Work
Table 5: Mixed effects modeling results from the expert user study. We report the standardized regression coefficient ($\beta$) and statistical significance (* $p<.05$, *** $p<.001$).
@@ 7 Conclusion
Limitations of ScholarEval. First, as a literature-grounded framework, ScholarEval relies heavily on the retrieved literature to evaluate research ideas. Hence, it is possible that it might misjudge a method’s effectiveness if it is not yet proven in the existing literature. However, the vast majority of ideas are not entirely novel and have some precedents in the literature (Wang et al., 2017). Second, as we mention in Appendix G, depending on the choice of the model backbone, running an idea evaluation using ScholarEval can take around 12 min and cost up to \$3. Although somewhat high, this remains on par with other literature search workflows (Skarlinski et al., 2024; Bragg et al., 2025). Future work may explore more lightweight versions of ScholarEval to enhance its usability to evaluate large batches of ideas.
@@
Limitations of Expert User Study. Some measures, such as those related to Focus, are inherently subjective to the user, and could be influenced by stylistic factors rather than actual improved quality. Additionally, resarch idea quality is a source of variance, as participants were only allowed to submit unique research ideas to a given system. However, we collect a self-report of research idea detail, which is not statistically significantly different according to a Welch’s t-test. Additionally, the overall scores of our results may be inflated by our expert sample, who report a 7.2/10 on AI use for research. Our results may be different in a sample who use AI less often. This does not affect our relative improvement over o4-mini-deep-research, as we show concrete evidence that the study was adequately blinded (see Appendix H). Future work should consider the barriers to adoption for automated research idea evaluation, especially for those who are less keen on using AI for their research cycle. Another strong direction would study how use of ScholarEval impacts short and long-term research success.
Ethical Considerations
ScholarIdeas is collected from papers on Openreview and eLife licensed under a Creative Commons Attribution (CC BY) license, which allows reuse with attribution. We provide the full list of papers used to create ScholarIdeas in Appendix D.
@@
ScholarIdeas is collected from papers on Openreview and eLife licensed under a Creative Commons Attribution (CC BY) license, which allows reuse with attribution. We provide the full list of papers used to create ScholarIdeas in Appendix D.
Acknowledgments
@@
Appendix A Reproducibility statement
To ensure the reproducibility of our work, we provide full prompts and implementation details of ScholarEval (Appendix B), the detailed process of constructing ScholarIdeas (Appendix D), full information about our evaluation (Appendix E), experimental setup (Appendix F), and expert user study (Appendix H). Upon acceptance, we will also openly release our code and dataset.
Appendix B ScholarEval details and prompts
@@
B.1 Soundness Details
B.1.1 Additional Implementation Details
@@
TL;DR Summary. Since the method-level soundness review can be lengthy, we also include a summarization step where we prompt an LLM to generate a TL;DR summary highlighting the top three most important strengths, weaknesses, and suggestions to address. In our ScholarEval user interface (Appendix C), both this summary and the method-level evaluation are shown to the user, with the latter being expandable.
@@ B.1.2 Prompts
Methods Extraction You are an expert research assistant. You are skilled at reading research ideas and identifying the methods that are being proposed to solve the research problem. Methods can be planned system designs, experiments, human studies, analyses, ablations, etc. Given a research idea, you should extract all methods as a Python list, such that each method is a separate item in the list. Each item should be a word-for-word copy of a method, along with a short synthesis that grounds the method in the context of the overall research idea. The extracted methods should be interpretable on their own. Ensure that the methods you extract address different aspects of the research idea and are non-redundant. The method list you return should be ranked by importance to the research idea, with the most important methods first. [start research idea] {research_idea} [end research idea] Please output a parseable Python block as follows: ‘‘‘python plans = ["context + method", ...] ‘‘‘
@@
Query Generation You are an expert research assistant. Given a method (i.e., one approach that researchers are adopting to execute their idea) extracted from a research idea, please construct a singular query that will be used to search for paper snippets using the Semantic Scholar API. Use JSON format with 70 words or less per query. Do not include any text in the query besides the query itself. Do not include text like "semantic search query about …" or "papers related to …". Just the actual query text. No operators such as AND, OR should be used. Just a query in natural language that is relevant to the method. [start extracted method] {clean_method} [end extracted method] In case the method does not have enough context to construct an effective snippet search query, you can use the research idea to understand the overall research direction and inject useful context. [start research idea] {research_idea} [end research idea] Please output a parseable JSON block as follows, being especially careful to use the correct number of escape characters: ‘‘‘json { "query": "Your search query here (IN 70 WORDS OR LESS)" } ‘‘‘
@@
Method and Result Synthesis You are an expert research assistant knowledgeable in many domains. You are extremely critical and observant, and do not overgeneralize findings. You are given a proposed research method and the methods/results section from a paper. [start paper] {paper_text} [end paper] [start proposed research method] {proposed_method} [end proposed research method] To further understand the scope of the proposed research method, here is the entire research idea that it is extracted from — a method is a single approach that researchers are adopting to execute their research idea: [start research idea] {research_idea} [end research idea] For any method in the paper that is related to the proposed research method and the overall research idea, please summarize the method used in the paper, report the experimental outcome from using the method, and provide some context for experimental conditions. Do not use any in-text citations. Ensure that the method, results, and context you provide are specific and detailed, and that they mention how it relates to the proposed method and research idea. If the proposed research method does not relate to any methods in the paper, please return an empty dictionary. Strictly follow the output format displayed below. JSON formatting requirements: - Must be a complete, valid JSON object - Start with an open bracket and end with closed bracket - No trailing commas after the last property - Validate JSON structure before output ‘‘‘json { "method": "Description of experimental approach including: algorithm/technique, datasets/inputs, computational resources, experimentation details, and evaluation setup, and metrics/instruments used, etc.", "results": "Quantitative outcomes with specific values, comparisons to baselines, statistical significance where applicable", "context": "Key experimental conditions: dataset/population size, hardware/system/instrument specs, hyperparameters, or other domain-specific constraints that affect reproducibility" } ‘‘‘
@@
Method-Level Soundness Review Synthesis You are an expert research assistant knowledgeable in many domains. You are extremely critical and observant, and do not overgeneralize findings. You are given a proposed research method and a list of related work. Your objective is to create a meta-review of the related work in the context of the proposed research method. Point out any evidence that supports or contradicts the proposed method. Make sure to contrast the related work as a series of iterative scientific work, where newer work can provide evidence that supports or contradicts older work. It is important that the meta-review you generate always ties back to the original research idea. Judge the support and contradictions as well as suggested actions for each method within the general context of the research idea to ensure that your review is highly relevant and precise. Be granular, making sure to reference specific details such as: algorithm/technique, datasets/inputs/population, computational resources, statistical methods, implementation details, and evaluation setup, and metrics/instruments used, etc. quantitative outcomes, comparisons to baselines, statistical significance dataset/population size, hardware/system specs, hyperparameters, or other domain-specific constraints that affect reproducibility It is important that for each method-level meta-review that you generate, your review of the support and contradictions should be ordered starting from strongest evidence of support/contradiction to the weakest. Likewise, the suggested actions should be ordered from most important to least important. This does not mean that you will generate these as bullet points, but rather detailed, coherent paragraphs that are logically ordered. It is required to copy the in-text citations with their links in Markdown format [(author, YYYY-MM)](link) when referring to related work. Related Work: {related_work} [start proposed research method] {pm} [end proposed research method] [start research idea] {rp} [end research idea] Output format: Please output a parseable JSON block as follows: ‘‘‘json { "support": "evidence that supports the proposed method", "contradictions": "evidence that contradicts the proposed method", "suggested_action": "how can the proposed method be improved based on the related work", "soundness_score": "int score 0 to 10 based on the evidence for and against the proposed method" } ‘‘‘
B.2 Contribution Details
B.2.1 Additional Implementation Details